A 12 pole, 3-phase alternator drives at a speed of 500 r.p.m supplies power to an 8 pole, three-phase induction motor. If the slip of the motor at full load speed is 4 %, calculate the full load of the motor. 


1
Expert's answer
2021-08-30T01:54:17-0400

 "N_s=\dfrac{120f}{p}\\
500=\dfrac{120f}{12}\\
f=50Hz\\
Speed\ of\ induction\ motor\\
N_s=\dfrac{120\times {50}}{8}=750rpm\\
N_r=N_s(1-s)\\
N_r=750(1-0.04)=720rpm\\"